# Multi-Modal Multi-Task Federated Foundation Models for Next-Generation Extended Reality Systems: Towards Privacy-Preserving Distributed Intelligence in AR/VR/MR

### v3

@@ Abstract
Extended reality (XR) systems, which consist of virtual reality (VR), augmented reality (AR), and mixed reality (XR), offer a transformative interface for immersive, multi-modal, and embodied human-computer interaction. In this paper, we envision that multi-modal multi-task (M3T) federated foundation models (FedFMs) can offer transformative capabilities for XR systems through integrating the representational strength of M3T foundation models (FMs) with the privacy-preserving model training principles of federated learning (FL). We present a modular architecture for FedFMs, which entails different coordination paradigms for model training and aggregations. Central to our vision is the codification of XR challenges that affect the implementation of FedFMs under the SHIFT dimensions: (1) Sensor and modality diversity, (2) Hardware heterogeneity and system-level constraints, (3) Interactivity and embodied personalization, (4) Functional/task variability, and (5) Temporality and environmental variability. We illustrate the manifestation of these dimensions across a set of emerging and anticipated applications of XR systems. Finally, we propose evaluation metrics, dataset requirements, and design tradeoffs necessary for the development of resource-aware FedFMs in XR. This perspective aims to chart the technical and conceptual foundations for context-aware privacy-preserving intelligence in the next generation of XR systems.
@@ Introduction
It is worth mentioning that FedFMs and M3T FedFMs are becoming increasingly practical, as evident in other fields [28, 29, 30], due to recent advances in resource-efficient local FM training/adaptation techniques. Further, in the AI/ML community, there is high interest in advancing parameter-efficient training/fine-tuning techniques, such as low rank adaptation (LoRA) and adapter modules, making the operations of M3T FedFMs over resource-constrained wireless devices increasingly practical [31].
@@
M3T FedFMs: Model Architecture and Network Orchestration
@@
1. Modality Encoders:
M3T FMs use dedicated modality encoders that convert raw sensory inputs into standardized intermediate embeddings. These encoders operate prior to fusion, serving as the entry point for the model backbone.
@@ 1. Modality Encoders:
5. FL Coordination Paradigm: Following FL principle, in M3T FedFM, each device performs local training over its dataset to update its local M3T FM modules/parameters; however, this process usually does not entail the training of the entire M3T FM, as the burden of such training exceeds the capabilities of resource-constrained devices. Instead, the local training is often performed in the form of prompt-tuning, learning the adapter parameters (e.g., via LoRA-based methods), or updating relevant M3T FM components, such as local experts. In Table 4, we provide a pairwise comparison between the techniques that can be used during the local training, identifying their unique advantages and disadvantages.
@@
In the following, we highlight a range of scenarios where M3T FMs have the potential to redefine how XR systems perceive, adapt, and interact, spanning domains such as healthcare, education, manufacturing, and entertainment. These scenarios, schematically visualized in Fig. 4, represent a blend of evolved versions of existing XR deployments and forward-looking use cases that push the boundaries of what is possible. Notably, with the recent wave of on-device FM integration into smartphones, particularly LLMs [42], it is increasingly clear that no major player in the XR space, including tech giants such as Apple, Google, and Meta can afford to ignore the paradigm shift that the recently emerged M3T FMs, and by extension, M3T FedFMs, are poised to bring to this field.
@@
In such settings, M3T FMs can enable truly intention-driven experiences, an emerging paradigm toward which both industry and academia are increasingly converging [47]. For instance, a user might simply think about a command on opening an application, and the interface seamlessly responds without requiring any spoken or gestural command. In the medical domain, this capability holds transformative potential [48]. For example, a patient with motor impairments might imagine moving a paralyzed limb. Subsequently, the M3T FM, trained to process multi-modal embodied inputs (e.g., neural signals, visual context, and spatial mapping), can interpret this mental intent and translate it into the movement of a virtual limb within the XR environment [49].
@@
In these systems, M3T FMs can lead to unprecedented advancements. For example, a companion powered by an M3T FM may detect a gradual decline in the user’s emotions, moods, or feelings (e.g., based on combined analysis of speech, gaze, and posture), and preemptively offer mental health resources. Also, by leveraging contextual cues such as eye-gaze and user behavior, AI companions can provide suggestions in terms of overlays to improve the user’s productivity [53].
@@
With modern XR headsets equipped with sensors for gaze tracking, pupil dilation, facial micro-expressions, and vocal stress detection, emotion-aware gaming can transition from a futuristic concept to an emerging reality, the promise of which has triggered a major interest in the gaming industry [54]. XR devices can continuously collect multi-modal embodied signals that reflect the player’s internal/emotional state in real-time. Building on this capability, adaptive storytelling is emerging as a new frontier in emotion-aware gaming, where narratives evolve not only in response to a player’s actions, but also in response to their physiological states [55, 56]. This enhances the emotional depth that XR games can offer, fostering a more profound and personalized connection between players and virtual environments. Notably, significant efforts have been directed toward dynamically eliciting/triggering complex emotions such as fear, creating deeply immersive and psychologically resonant gaming experiences [57, 58].
@@
Why in the Above (and All Comparable) Scenarios M3T FMs Must be Accompanied by FL? A common thread across all the scenarios discussed above is the geo-distributed nature of data generation. XR systems deployed in homes, hospitals, factories, classrooms, and entertainment spaces continuously collect embodied, multi-modal data, which is generated/collected locally on XR devices. Such a geo-distributed nature of data generation poses a major challenge to the integration of M3T FMs into XR scenarios: traditional centralized M3T FM training paradigms require transferring/pooling data from all XR devices into a centralized server to train a globally shared model. Yet, in the context of XR, the centralization of XR data is impeded by three key limitations: Privacy: XR data inherently includes deeply personal and behavioral information (e.g., emotional states, cognitive intentions, and biometric patterns) that are not suitable for off-device transmission or cloud storage. Bandwidth and Latency: Continuously uploading high-dimensional sensory data from distributed XR devices imposes unsustainable demands on wireless infrastructure, introducing unacceptable delays in real-time experiences. Personalization: Centralized training typically produces generic models that fail to reflect local variations in user behavior and context – critical for applications such as neuroadaptive interfaces or emotion-aware storytelling. $\star$ Leveraging the distributed model training capabilities of FL, M3T FedFMs offer a compelling alternative that can address these challenges, making them a major component of next-generation intelligent XR systems.
@@
SHIFT Dimensions and their Manifestation in XR Ecosystems
The implementation of FedFMs in the XR domain, despite their high potential, faces a variety of nuanced challenges that must be carefully characterized in order to guide future research, system design, and deployment strategies. Motivated by the need to isolate and understand the most relevant obstacles, we categorize the core challenges associated with training, adapting, and personalizing FedFMs in XR systems under SHIFT dimensions: (1) Sensor and modality diversity, (2) Hardware heterogeneity and system-level constraints, (3) Interactivity and embodied personalization, (4) Functional/task variability, and (5) Temporality and environmental variability. It is worth mentioning that these challenges/dimensions are complementary to those of conventional FL (e.g., data heterogeneity [59]).
Below, we elaborate on each of SHIFT dimensions, highlighting how it imposes constraints that can disrupt standard FedFM learning assumptions and necessitates a rethinking of how FedFMs are designed, trained, and deployed in XR environments.
@@
XR devices may differ vastly in the types and combinations of sensors they possess, ranging from lightweight AR glasses with vision-only input to immersive VR headsets equipped with eye-tracking, haptic feedback, spatial audio, and depth sensing. This sensor heterogeneity leads to non-overlapping feature spaces across devices. Moreover, modalities may dynamically appear or disappear within an XR session (e.g., a muted microphone, disabled haptics, or environmental occlusion). Addressing these characteristics in conjunction requires enabling XR devices with incomplete modality views and temporal modality dropouts and emergence to still contribute meaningfully to FedFMs, which are not conventionally designed to handle such characteristics.
@@
Metric Description How to Measure SHIFT Dimensions
@@
Inference Latency Time between the reception of sensing inputs and model output generation; critical for responsiveness. Log time between input event (e.g., gesture, gaze) and output generation. H  –  I  –  T
@@
Personalization Effectiveness Performance gain after local personalized (e.g., using prompt tuning or local training) vs. generic global model. Compare task accuracy or user satisfaction rate before and after personalization. I  –  F
@@
Personalization Necessity Rate Proportion of users or tasks where global FedFM meets performance threshold, avoiding need for adaptation. Count users/tasks for which the global model exceeds a baseline accuracy (e.g., 90%) without fine-tuning. H  –  I  –  F
@@
Cross-Context Generalization The model’s ability to generalize effectively across previously unseen users, environments, or tasks. Test model on data from new users or contexts not seen during training. S  –  F  –  T
@@
Task-Specific Continuity Retaining performance on previously learned tasks during adaptation on the new tasks. Evaluate drop in earlier task accuracy post new-task fine-tuning (i.e., model forgetting measure under continual learning). F  –  T
@@
Submodule Activation Efficiency Average number of active FedFM modules per task; reflects model modularity. Count number of modules activated per task or per user interaction. H  –  F
@@
Model Storage Footprint Total model size deployed on XR device, including adapters or quantized formats. Sum memory usage in MB for all active modules on device. H
@@
Communication Overhead Amount of data transferred per aggregation round or XR session. Monitor the bits exchanged during global aggregation rounds or XR sessions. H  –  T
@@
Battery Impact / Energy Use Device power consumption while running/adapting FedFM. Measure power draw in mW or energy used (Joules) via onboard or external sensors. H  –  T
@@
User-Reported QoE Subjective user-perceived quality of experience, focusing on responsiveness and interaction fluidity. Collect Likert-scale ratings, qualitative feedback, or lag reports through user surveys. I  –  T
@@
Model-Inferred QoE System-side estimation of perceived interaction quality without explicit user input. Monitor implicit behavioral signals such as gaze jitter, task repetition, latency-induced hesitation, or input aborts. I  –  T
@@
On-Device Training Time Duration required for local model adaptation, tuning, or fine-tuning directly on the XR device. Record duration of local fine-tuning routine on device per adaptation event. H  –  I  –  T
@@
Privacy Leakage Risk Potential for data leakage through observation of model updates or representations during global aggregations. Run membership inference attacks, gradient inversion, or compute differential privacy bounds. S
@@
XR environments are not passive; instead, they are inherently interactive, responding to user’s motion, attention, speech, and preference in real-time. This interactivity, combined with the embodied nature of XR (i.e., user movement, posture, and environment are part of the interaction loop), introduces new challenges for FedFM personalization: FedFMs must adapt continuously in XR devices, rather than post-hoc. Nevertheless, conventional FedFM personalization techniques, such as local fine-tuning, personalized output heads or adapters, are often resource-demanding or not quick/agile enough for XR scenarios, where real-time feedback and low-latency inference are critical.
@@
In a single session/time-window of XR interactions, the user’s tasks may vary from identifying objects and following spatial navigation cues, to engaging in conversations with a virtual assistant and responding to real-time gesture-based commands. These tasks differ not only in output format but also in temporal scale (e.g., real-time vs. near-real-time), relevant modality, and computational demand. This intra-user multi-tasking in XR is further compounded by inter-user and inter-application variability: a user in a manufacturing setting may need real-time error detection and instructions, while a student in an educational XR system might engage in immersive simulations and voice-based tutoring. This breadth of functionality, which we refer to as functional variability, breaks the common assumption in conventional FedFMs that devices often share a fixed set of downstream tasks or training objectives.
@@
  • Robust Multimodal Fusion under Missing Inputs: Develop FedFM models that can operate effectively with partial, missing, or dynamically changing sensory streams by leveraging redundancy across modalities.
@@
  • Composable User-Specific Modules: Optimize the selection and use of the plug-and-play components (e.g., adapters or task heads) that specialize rapidly to a user’s context without modifying the global model backbone.

### v4

@@ Abstract
Extended reality (XR) systems, encompassing virtual reality (VR), augmented reality (AR), and mixed reality (MR), offer a transformative interface for immersive, multi-modal, and embodied human-computer interaction. In this paper, we envision that multi-modal multi-task (M3T) federated foundation models (FedFMs) can offer transformative capabilities for XR systems through integrating the representational strength of M3T foundation models (FMs) with the privacy-preserving and personalized model training principles of federated learning (FL). To this end, we first present the modular architecture of FedFMs, which entails different coordination paradigms for model training and aggregations. Afterward, we codify XR challenges that affect the implementation of FedFMs under the SHIFT dimensions: (1) Sensor and modality diversity, (2) Hardware heterogeneity and system-level constraints, (3) Interactivity and embodied personalization, (4) Functional/task variability, and (5) Temporality and environmental variability. We then illustrate the manifestation of these dimensions across a set of emerging and anticipated applications of XR systems. Finally, we propose evaluation metrics, dataset requirements, and design tradeoffs necessary for the development of resource-efficient FedFMs in XR ecosystems.
@@ Introduction
Although training challenges and stringent inference requirements, especially in XR environments, pose potential hurdles for FedFMs, recent trends suggest a promising trajectory. FedFMs and M3T FedFMs are becoming increasingly practical, as evident in other fields [28, 29, 30], driven by recent advances in resource-efficient local FM training/adaptation techniques and the emergence of compact, high-efficiency models such as Google DeepMind’s Gemma 3n [31], Meta Llama 3.2 [32], and Apple Intelligence [33], which are specifically engineered for real-time, on-device multi-modal inference. Further, in the AI/ML community, there is high interest in advancing parameter-efficient training/fine-tuning techniques, such as low rank adaptation (LoRA) and adapter modules, making the operations of M3T FedFMs over resource-constrained wireless devices increasingly practical [34].
@@
M3T FedFMs: Model Architecture, Network Orchestration, and Training Strategy
@@
1. Modality Encoders: M3T FMs use dedicated modality encoders that convert raw sensory inputs into standardized intermediate embeddings. These encoders operate prior to fusion, serving as the entry point for the model backbone.
@@
5. FL Coordination Paradigm: Following FL principle, in M3T FedFM, each device performs local training over its dataset to update its local M3T FM modules/parameters; however, this process usually does not entail the training of the entire M3T FM, as the burden of such training exceeds the capabilities of resource-constrained devices. Instead, the local training is often performed in the form of prompt-tuning, learning the adapter parameters (e.g., via LoRA-based methods [41]), or updating relevant M3T FM components, such as local experts. In Table 4, we provide a pairwise comparison between the techniques that can be used during the local training, identifying their unique advantages and disadvantages.
@@
6. Training Strategies of M3T FedFMs: Effective training of M3T FedFMs demands a fundamental reassessment of conventional FL techniques, given the architectural and functional shifts introduced by M3T models. Techniques such as model distillation, quantization, sparsification, intelligent client selection, and parameter-efficient fine-tuning (PEFT), such as LoRA, have been predominantly developed for monolithic, task-specific deep neural networks or LLMs, where their extensions to the modular and multi-modal multi-task nature of emerging M3T FedFMs remain highly underexplored [46, 39]. A promising direction lies in adapting and extending these techniques to align with the modular structure of M3T models. Specifically, their modular design, featuring expert modules, task-specific heads, and lightweight adapters, enables selective/module-wise training and fine-tuning on client devices, followed by modular-level aggregations, which calls for the revisiting of these existing techniques. For instance, model distillation (traditionally used to compress a large model into a smaller one) can evolve into a modality-aware distillation process that preserves knowledge across different sensor modalities and tasks without collapsing their distinct representations. Similarly, quantization and sparsification can be applied selectively to different modules (e.g., encoders vs. task heads), balancing compression with model fidelity. Also, intelligent client selection (typically relying on system metrics such as availability or data quality) can be expanded to incorporate modality-task coverage, prioritizing clients that possess underrepresented sensor combinations or task labels. Finally, PEFT methods such as LoRA or adapters (originally designed for sequential fine-tuning of LLMs) should be reconfigured to support multi-branch, modular updates (i.e., updates targeting specific components of the model such as modality-specific encoders, task heads, or expert modules), allowing clients to locally fine-tune only relevant modules and transmit them for federated aggregation. Taking this one step further, when deploying M3T FedFMs in XR systems, these revamped techniques must also account for the unique interaction, resource, and environment characteristics of XR, which we will later elaborate on through the SHIFT dimensions in the following sections.
@@
In the following, we highlight a range of scenarios where M3T FMs have the potential to redefine how XR systems perceive, adapt, and interact, spanning domains such as healthcare, education, manufacturing, and entertainment. These scenarios, schematically visualized in Fig. 4, represent a blend of evolved versions of existing XR deployments and forward-looking use cases that push the boundaries of what is possible. Notably, with the recent wave of on-device FM integration into smartphones  [31, 32, 33], it is increasingly clear that no major player in the XR space, including tech giants such as Apple, Google, and Meta can afford to ignore the paradigm shift that the recently emerged M3T FMs, and by extension, M3T FedFMs, are poised to bring to this field.
@@
In such settings, M3T FMs can enable truly intention-driven experiences, an emerging paradigm toward which both industry and academia are increasingly converging [51]. For instance, a user might simply think about a command on opening an application, and the interface seamlessly responds without requiring any spoken or gestural command. In the medical domain, this capability holds transformative potential [52, 53]. For example, a patient with motor impairments might imagine moving a paralyzed limb. Subsequently, the M3T FM, trained to process multi-modal embodied inputs (e.g., neural signals, visual context, and spatial mapping), can interpret this mental intent and translate it into the movement of a virtual limb within the XR environment [54].
@@
In these systems, M3T FMs can lead to unprecedented advancements. For example, a companion powered by an M3T FM may detect a gradual decline in the user’s emotions, moods, or feelings (e.g., based on combined analysis of speech, gaze, and posture), and preemptively offer mental health resources [58, 59]. Also, by leveraging contextual cues such as eye-gaze and user behavior, AI companions can provide suggestions in terms of overlays to improve the user’s productivity [60].
@@
With modern XR headsets equipped with sensors for gaze tracking, pupil dilation, facial micro-expressions, and vocal stress detection, emotion-aware gaming can transition from a futuristic concept to an emerging reality, the promise of which has triggered a major interest in the gaming industry [61]. XR devices can continuously collect multi-modal embodied signals that reflect the player’s internal/emotional state in real-time [62]. Building on this capability, adaptive storytelling is emerging as a new frontier in emotion-aware gaming, where narratives evolve not only in response to a player’s actions, but also in response to their physiological states [63, 64]. This enhances the emotional depth that XR games can offer, fostering a more profound and personalized connection between players and virtual environments. Notably, significant efforts have been directed toward dynamically eliciting/triggering complex emotions such as fear, creating deeply immersive and psychologically resonant gaming experiences [65, 66].
@@
Why in the Above (and All Comparable) Scenarios M3T FMs Must be Accompanied by FL? A common thread across all the scenarios discussed above is the geo-distributed nature of data generation. XR systems deployed in homes, hospitals, factories, classrooms, and entertainment spaces continuously collect embodied, multi-modal data, which is generated/collected locally on XR devices. Such a geo-distributed nature of data generation poses a major challenge to the integration of M3T FMs into XR scenarios: traditional centralized M3T FM training paradigms require transferring/pooling data from all XR devices into a centralized server to train a globally shared model. Yet, in the context of XR, the centralization of XR data is impeded by three key limitations: Privacy: XR data inherently includes deeply personal and behavioral information (e.g., emotional states, cognitive intentions, and biometric patterns) that are not suitable for off-device transmission or cloud storage. Bandwidth and Latency: Continuously uploading high-dimensional sensory data from distributed XR devices imposes unsustainable demands on wireless infrastructure, introducing unacceptable delays in real-time experiences. Personalization: Centralized training typically produces generic models that fail to reflect local variations in user behavior and context – critical for applications such as neuroadaptive interfaces or emotion-aware storytelling. $\star$ Leveraging the distributed model training capabilities of FL, M3T FedFMs offer a compelling alternative that can address these challenges, making them a major component of next-generation intelligent XR systems.Specifically, as elaborated previously and illustrated in Fig. 3, integrating FL into M3T FedFMs enables on-device model training/fine-tuning, ensuring that user data remains local, either on the originating XR device or within the user’s personal ecosystem of devices. This architecture prevents raw user data (e.g., emotional states, biometric signals) from being exposed to external entities as only model/gradients parameters are shared across the system. Moreover, by avoiding real-time data uploads and keeping voluminous sensory inputs local at user devices, this approach mitigates bandwidth and latency concerns. Finally, the inherent support for local model adaptation in FL allows M3T FedFMs to be tailored to individual users’ behaviors and contexts, thereby offering model personalization.
@@
Manifestations of SHIFT Dimensions: Why XR Represents a Unique Frontier for FedFMs?
Although the integration of FL into FedFMs helps address the core challenges of FMs in XR ecosystems (namely, challenges related to ‘Privacy’, ‘Bandwidth and Latency’, and ‘Personalization’ as discussed above), the practical implementation of FedFMs in XR environments introduces a range of nuanced challenges. While several of these challenges, such as hardware heterogeneity, communication cost, and heterogeneous/non-IID data distribution, are shared with other FedFMs settings, their manifestation in XR environments is uniquely intensified by the real-time, immersive, and multi-modal nature of user interaction. In particular, XR devices must locally run large FMs while handling diverse multi-modal inputs, often in dynamic physical and virtual environments. These systems operate under strict constraints on power, compute, memory, storage, and thermal capacity, while simultaneously requiring low-latency, personalized responses to support interactive, task-varying, and context-aware user experiences. These challenges must be carefully characterized to inform future research directions, system design, and deployment strategies. Motivated by the need to isolate and understand the most relevant obstacles, we categorize the core challenges associated with training, adapting, and personalizing FedFMs in XR systems under SHIFT dimensions: (1) Sensor and modality diversity, (2) Hardware heterogeneity and system-level constraints, (3) Interactivity and embodied personalization, (4) Functional/task variability, and (5) Temporality and environmental variability. It is worth mentioning that these challenges/dimensions are complementary to those of conventional FL (e.g., data heterogeneity [67]).
Below, we elaborate on each of SHIFT dimensions, highlighting how it imposes constraints that can disrupt standard FedFM learning assumptions and necessitate a rethinking of how FedFMs are designed, trained, and deployed in XR environments.
@@
XR devices may differ vastly in the types and combinations of sensors they possess, ranging from lightweight AR glasses with vision-only input to immersive VR headsets equipped with eye-tracking, haptic feedback, spatial audio, and depth sensing. This sensor heterogeneity results in non-overlapping feature spaces across devices. Additionally, input modalities may dynamically appear or disappear within an XR session: for instance, due to muted microphones, disabled haptics, or transient occlusions in the environment. If left unaddressed, this heterogeneity and temporal variability in sensor modalities can severely degrade the performance and convergence of FedFMs. Specifically, devices with incomplete or shifting modality views may contribute noisy, biased, or incompatible model updates to the global model, leading to poor generalization, slower convergence, or even divergence during collaborative model training. Therefore, it is critical to design FedFMs that can robustly incorporate updates from devices with incomplete modality views and temporally dynamic input streams, ensuring that all participating devices, regardless of their sensing capabilities or environmental conditions, can contribute meaningfully to the collaborative model training process.
@@
Metric Description How to Measure SHIFT Dimension Relevance
@@
Inference Latency Time between the reception of sensing inputs and model output generation; critical for responsiveness. Log time between input event (e.g., gesture, gaze) and output generation. Real-time interaction (I) on diverse XR hardware (H) under varying conditions (T).
@@
Personalization Effectiveness Performance gain after local personalized (e.g., using prompt tuning) vs. generic global model. Compare task accuracy or user satisfaction rate before and after personalization. User-specific interaction (I) by adapting to diverse task requirements (F) in XR interfaces.
@@
Personalization Necessity Rate Proportion of users or tasks where global FedFM meets threshold, avoiding need for adaptation. Count users/tasks for which the global model exceeds a baseline accuracy (e.g., 90%) without fine-tuning. Balancing hardware use (H) and user needs (I) across tasks (F) in XR systems.
@@
Cross-Context Generalization The model’s ability to generalize effectively across previously unseen users, environments, or tasks. Test model on data from new users or contexts not seen during training. Diverse sensor inputs (S), tasks (F), and dynamic XR environments (T).
@@
Task-Specific Continuity Retaining performance on previously learned tasks during adaptation on the new tasks. Measure accuracy drop on prior tasks after new-task fine-tuning (i.e., forgetting under continual learning). Performance consistency across evolving tasks (F) in dynamic XR settings (T).
@@
Submodule Activation Efficiency Average number of active FedFM modules per task; reflects model modularity. Count number of modules activated per task or per user interaction. Optimized resource use on varied XR hardware (H) for diverse tasks (F).
@@
Model Storage Footprint Total model size deployed on XR device, including adapters or quantized formats. Sum memory usage in MB for all active modules on device. Minimizing storage demands on resource-constrained XR hardware (H).
@@
Communication Overhead Amount of data transferred per aggregation round or XR session. Monitor the bits exchanged during global aggregation rounds or XR sessions. Data transfer for efficient operation on XR devices (H) in dynamic contexts (T).
@@
Battery Impact / Energy Use Device power consumption while running/adapting FedFM. Measure power draw in mW or energy used (Joules) via onboard or external sensors. Energy efficiency on XR hardware (H) for prolonged use in varied environments (T).
@@
User-Reported QoE Subjective user-perceived quality of experience, focusing on responsiveness and interaction fluidity. Collect Likert-scale ratings, qualitative feedback, or lag reports through user surveys. Immersive, responsive XR interactions (I) in dynamic settings (T).
@@
Model-Inferred QoE System-side estimation of perceived interaction quality without explicit user input. Monitor implicit behavioral signals like gaze jitter, task repetition, latency-induced hesitation, or input aborts. Seamless XR interaction (I) by detecting quality issues in real-time environments (T).
@@
On-Device Training Time Duration required for local model adaptation, tuning, or fine-tuning directly on the XR device. Record duration of local fine-tuning routine on device per adaptation event. Speed of local adaptation on XR hardware (H) for personalized interactions (I) in dynamic contexts (T).
@@
Privacy Leakage Risk Potential data leakage via observation of model updates or representations during global aggregations. Run membership inference attacks, gradient inversion, or compute differential privacy bounds. Protecting sensitive sensor data (S) in XR ecosystems with diverse inputs.
@@
XR environments are not passive; instead, they are inherently interactive, responding to user’s motion, attention, speech, and preference in real-time. This interactivity, coupled with the embodied nature of XR (where user movement, posture, and environment are integral to the interaction loop), introduces unique challenges for FedFM personalization [69]. Specifically, FedFMs must maintain generalizability while enabling continuous, low-latency, on-device adaptation. In this context, the collaborative training paradigm of FL, as integrated into M3T FedFMs, allows devices to aggregate locally trained/fine-tuned model components, thereby enhancing model generalizability compared to isolated, non-collaborative training. Additionally, the previously described fine-tuning techniques employed in M3T FMs can potentially enable effective model personalization. However, these fine-tuning strategies (e.g., prompt tuning, adding personalized task heads, or introducing adapters) may fall short in XR contexts, as they can be too resource-intensive or insufficiently responsive to support the real-time, low-latency demands of XR applications. Additionally, it remains unclear how frequently such local personalization should be triggered in practical XR scenarios, where user states and interactions evolve continuously in real-time.
@@
In a single session/time-window of XR interactions, the user’s tasks may vary from identifying objects and following spatial navigation cues, to engaging in conversations with a virtual assistant and responding to real-time gesture-based commands. These tasks differ not only in output format but also in temporal scale (e.g., real-time vs. near-real-time), relevant modality, and computational demand. This intra-user multi-tasking in XR is further compounded by inter-user and inter-application variability: a user in a manufacturing setting may need real-time error detection and instructions [70], while a student in an educational XR system might engage in immersive simulations and voice-based tutoring [71]. This breadth of functionality, which we refer to as functional variability, breaks the common assumption in conventional FedFMs that devices often share a fixed set of downstream tasks or training objectives.
@@
  • Robust Multimodal Fusion under Missing Inputs: Develop FedFM models that can operate effectively with partial, missing, or dynamically changing sensory streams by leveraging redundancy across modalities and data augmentation[82].
@@
  • Composable User-Specific Modules: Design and optimize XR-specific plug-and-play components (e.g., adapters, task heads) that can quickly specialize to a user’s context without altering the shared global backbone. These modules could also be sourced from other users who have encountered similar data or interaction patterns, though identifying and transferring such modules across users remains an open challenge.